### v2

# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

Abstract
The basic reproduction number ($R_{0}$) is an epidemiological metric that represents the average number of new infections caused by a single infectious individual in a completely susceptible population. The methodology for calculating this metric is well-defined for numerous model types, including, most prominently, Ordinary Differential Equations (ODEs). The basic reproduction number is used in disease modeling to predict the potential of an outbreak and the transmissibility of a disease, as well as by governments to inform public health interventions and resource allocation for controlling the spread of diseases. A Petri net (PN) is a directed bipartite graph where places, transitions, arcs, and the firing of the arcs determine the dynamic behavior of the system. Petri net models have been an increasingly used tool within the epidemiology community. However, a generalized method for calculating $R_{0}$ directly from PN models has not been established. Thus, in this paper, we present a general method for calculating $R_{0}$ for Petri nets. Additionally, we show how a computational method implementing the next-generation algorithm in ODE models can also be applied to Petri net models. We also provide multiple examples of how to use this approach to calculate $R_{0}$ for various SIR-type Petri net models.
keywords: Petri Nets, Reproduction Number, SIR model, Disease Spread, Next-Generation Method, Ordinary Differential Equations, Epidemiology, Modeling
1 Introduction
In the context of epidemiology, the basic reproduction number ($R_{0}$) refers to the expected number of new cases that one infected individual is likely to cause in a population where the entire population is susceptible [1]. As such, $R_{0}$ is a fundamental concept in mathematical epidemiology for all communicable diseases when it comes to Ordinary Differential Equation (ODE) based models. Intuitively, $R_{0}$, represents the expected number of secondary infections caused by a single infected individual in a fully susceptible population [1, 2]. An accurate calculation of $R_{0}$ is therefore crucial to understanding an infectious disease’s potential spread and informing public health interventions[3]. Analyzing $R_{0}$ can give insights into the future dynamics of a particular disease based on the parameters, subpopulations, and constants that make up any given $R_{0}$ value. Researchers calculate $R_{0}$ in models to assess transmission dynamics, compare scenarios, and evaluate the effectiveness of control measures such as vaccination or social behavior [3]. Mathematical treatments of $R_{0}$ via the ”next-generation matrix method” (NGM) have been outlined by Diekmann [2, 4, 5] and Van den Driessche [6]. Although other approaches exist for calculating $R_{0}$, the next-generation method is one of the most commonly used for ODE models[3] and so we focus on it here.
Petri nets provide a promising alternative to standard Ordinary Differential Equation models in epidemiology [7, 8, 9]. Petri net models are easily visualized as causal network diagrams, and their modular structure is possible to scale to large spatial models [10]. They can also integrate continuous time models of disease spread with discrete-time events such as policy changes or interventions. Numerous studies have explored the use of Petri nets to represent and analyze SIR models, highlighting their potential for modular and large scale modeling of disease dynamics [11, 12, 13, 14, 15, 16]. However, unlike the well-developed and widely applicable techniques for computing the basic reproduction number $R_{0}$ in ODE models, no comparably general method exists for deriving $R_{0}$ directly from Petri net representations. The inability to directly calculate $R_{0}$ for Petri nets limits the analytical utility of Petri nets and poses a significant barrier to the broader adoption for the modular construction of large scale, mechanistic models of infectious disease spread.
Applying the next-generation matrix method of calculating $R_{0}$ within the framework of Petri nets requires insights on how to bridge the gap between mechanistic model structure and formal spectral analysis. The contribution of this paper is introducing those insights in how the transition and transmission matrices are defined and characterizing $R_{0}$ in the context of Petri nets. Among the various formulations of the next-generation matrix method, the computational approach introduced by Diekmann et al. [2, 4, 5] is uniquely well-suited for implementation in Petri nets, as it explicitly constructs the next-generation matrix from compartmental transitions, which naturally aligns with the token flow and transition structure of Petri nets. This compatibility allows for a direct mapping from Petri net dynamics to the linearized infection and transition matrices required for $R_{0}$ calculation.
In this work, we demonstrate the applicability and generality of this approach across a range of epidemiological models, including classical structures like the SIRS and SEEIR models, a compartmental COVID-19 model with various infected compartments, and a model with a nonlinear infection term. These examples collectively illustrate how the Diekmann formulation adapted to Petri nets enables rigorous and automated computation of $R_{0}$ in Petri nets, supporting their use across a broader spectrum of infectious disease modeling.
The remainder of the paper is structured as follows. In Section 2, we introduce the Petri net formalism and two ways of using PNs to implement Susceptible-Infectious-Recovered (SIR) models. One approach uses deterministic, discrete-time PNs with variable arc weights and the other uses stochastic, continuous-time PNs with constant arc weights representing the mass action of interactions among individuals in the compartments. Section 3 reviews Diekmann’s approach to calculating $R_{0}$ and shows how it applies to a range of example model types, including non-linear models and spatial patch systems. Section 4 presents a numerical verification of our approach by comparing the calculated PN $R_{0}$ values to the corresponding ODE $R_{0}$ values.
2 A Brief Overview of Petri Nets
A Petri net graph, or Petri net structure, is a weighted bipartite graph [17] defined as $n$-tuple $(P,T,A,w,x)$ where:
  • $P$ is the finite set of places (one type of vertex in the graph).
  • $T$ is the finite set of transitions (the other type of vertex in the graph).
  • $A$ is the set of arcs (edges) from places to transitions and from transitions to places in the graph $A\subset(P\times T)\cup(T\times P)$.
  • $M_{0}$ is the initial state (also known as marking), $M_{0}=[m_{1},m_{2},\ldots,m_{n}]$, where $m_{i}$ is the number of tokens in place $p_{i}$
  • $w$ : A → {1, 2, 3, . . .} is the weight function on the arcs.
  • $x$ is a marking of the set of places P; $x=[x(p_{1}),x(p_{2}),...,x(p_{n})$] $\in N^{n}$ is the row vector associated with $x$.
Tokens are assigned to places, with the initial assignment being the initial marking. The number of tokens assigned to a place is an arbitrary non-negative integer but does not necessarily have an upper bound. A transition $t_{j}\in T$ in a Petri net is said to be enabled if $x(p_{i})\geq w(p_{i},t_{j})$ for all $p_{i}\in I(t_{j})$, where $I(t_{j})$ is the set of input arcs from places to $t_{j}$. This allows us to define the state transition function, $f:{\mathbb{N}}^{n}\times T\to{\mathbb{N}}^{n}$, of Petri net $(P,T,A,w,x)$ is defined for transition $t_{j}\in T$ if and only if $x(p_{i})\geq w(p_{i},t_{j})$ for all $p_{i}\in I(t_{j})$. If $f(x,t_{j})$ is defined, then we set $x^{\prime}=f(x,t_{j})$, where $x^{\prime}(p_{i})=x(p_{i})-w(p_{i},t_{j})+w(t_{j},p_{i}),i=1,...,n$. In simple terms, a transition is enabled if the number of tokens in all places connected to that transition via an incoming arc is greater than or equal to the arc weight for the respective arc connected to the transition.
An important point to note is that there may be multiple possible PN representations for a given ODE model because there are choices in how to implement the PN model. The most common mapping method follows a method of mapping ODEs to Stochastic Petri nets (SPNs) as laid out by Baez and Biamonte [18]. An alternative method is to map the ODE to a Variable Arc Weight Petri net (VAPN) [7]. Variable arc weights Petri nets can have arc weights that depend on parameters, the number of tokens in places at specific times, and time. This method is deterministic and invertible, allowing for potentially lower computation in parameter fitting and simulations, but it is novel and not yet widely used.
3 Next-Generation Matrix For Petri Nets
We first review the next-generation method for ODEs before showing how we can apply it to PNs.
3.1 Next-Generation Matrix for Ordinary Differential Equations
The method for calculating $R_{0}$ via the Next-Generation Matrix (NGM) laid out by Diekmann et al. [2, 4, 5] defines $R_{0}$ as the spectral radius of the next-generation matrix. Specifically, two matrices (transmission and transition matrices) are defined to find $R_{0}$, starting with $T$, which is defined as the transmission matrix. The transmission matrix outlines the production of new infections. Where elements within the matrices describe how people move from non-infected compartments to infected compartments. Next, they define $\Sigma$ as the transition matrix, which describes state changes due to individuals moving to other infected compartments, dying, or becoming immune. To then find $R_{0}$ we take the dominant eigenvalue of $(-T\Sigma^{-1})$ yielding $R_{0}=\varphi(-T\Sigma^{-1})$.
Alternatively, van den Driessche and Watmough [6] outline a similar definition of $R_{0}$. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of individuals in each compartment. Let $n$ be the number of compartments with the first $m$ compartments being the infected compartments. We define $X_{s}$ as the set of all disease free states such that $X_{s}=\{x\geq 0|x_{i}=0,i=1,...,m\}$. Then we define the transmission matrix, matrix $\mathcal{F}_{i}(x)$ to be the rate of appearance of new infections in compartment $i$. We use two matrices $\mathcal{V}^{+}_{i}(x)$ and $\mathcal{V}^{-}_{i}(x)$ to describe the rate of transfer of individuals into compartment $i$ by all other means and the rate of transfer of individuals out of compartment $i$, respectively. From $\mathcal{V}^{+}_{i}(x)$ and $\mathcal{V}^{-}_{i}(x)$, we get the transition matrix, $\mathcal{V}_{i}=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x).$ The disease model itself has non-negative initial conditions with the system of equations following
$\dot{x_{i}}$ $=$ $f_{i}(x)=\mathcal{F}_{i}(x)-\mathcal{V}_{i}(x),i=1,...n.$ (1)
While some of these assumptions are trivial when modeled in Petri nets, others, like derivations, need to be inspected more closely. For reference, we state the five mathematical assumptions here exactly as they are given in van den Driessche and Watmough [6] followed by brief comments. We will cover these assumptions for Petri nets in the following Section 3.
  1. (1)
    If $x\geq 0$, then $\mathcal{F}_{i},\mathcal{V}^{+}_{i},\mathcal{V}^{-}_{i}\geq 0$ for $i=1,…,n$. Each subsequent function $\dot{x_{i}}$ represents a transfer of individuals, thus they are all non negative.
  2. (2)
    If $x_{i}=0$ then $\mathcal{V}^{-}_{i}=0$. In particular, if $x\in X_{s}$ , then $\mathcal{V}^{-}_{i}=0$ for $i=1,…,m$. If a compartment is empty, then there can be no transfer of individuals out of the compartment.
  3. (3)
    $\mathcal{F}_{i}=0$ if $i>m$. No infectious population enters a non-infectious compartment.
  4. (4)
    If $x\in X_{s}$ then $\mathcal{F}_{i}(x)=0$ and $\mathcal{V}^{+}_{i}(x)=0$ for $i=1,…,m$. This simply states that if the population is disease-free, it stays disease-free. Meaning no, density-independent infections population immigrates in. This becomes important when dealing with patch models, as seen in Example 4.6.1.
  5. (5)
    If $\mathcal{F}(x)$ is set to zero, then all eigenvalues of $Df(x_{0})$ have negative real parts, where $Df(x_{0})$ is the derivative $[\frac{\partial f_{i}}{\partial x_{j}}]$ evaluated at the disease free equilibrium (DFE), $x_{0}$.
These assumptions allow for the proof of Lemma 1 from van den Driessche and Watmough [6] and using Lemma 1, van den Driessche and Watmough go on to prove $\varphi(FV^{-1})=R_{0}$.
3.2 Next-Generation Matrix for Petri Nets
Since many of the papers we draw on for examples below use the van den Driessche and Watmough notation, we will lay out the Next-Generation Matrix for Petri nets (NGMPN) following similar notation. To help in visualizing some of the concepts outlined in this process Figure 1 will be used as a basic VAPN model with non-infected places ($N_{1},N_{2}$) and infected places ($I_{1},I_{2},I_{3}$). Examples using SPNs will be given in the Case Studies section under various cases.
Figure 1: This VAPN is used as a guide to help visualize some of the concepts described in the Next-Generation Matrix for Petri Nets formulation. The places of this VAPN can be labeled $(P_{1},P_{2},P_{3},P_{4},P_{5})$ or equivalently $(N_{1},I_{1},I_{2},I_{3},N_{2})$, where the second label gives more description on the infectious state of the place.
First, think of places in Petri nets the same way that you might think about compartments in ODEs. Then we assume we have a Petri net model for a disease that fits the definition of VAPN or SPN laid out in Section 2. Assume that each transition can fire once per time step. Define the total amount infected places in the Petri net model as $T$. Thus, the infected places are defined as $I_{1},I_{2},...,I_{T}$. As a note, we define all places as $P$ and they are labeled $P_{1},...,P_{L}$, where $L$ is total number of places and $T\leq L$. We define $t_{N}$ as a transition with an input arc coming from a non-infected place and an output arc going to an infected place. With $w(t_{N_{j,...,k}},I_{i})$ being the sum of all arcs weights $j,...,k$ from arcs going from non-infected places going to an infected place $I_{i}$ for VAPN or the net sum of all rate equations $j,...,k$ for arcs from transitions with input arcs from non-infected places and output arcs going to infected places $I_{i}$ for SPN. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of tokens in each place, being the Petri net equivalent of population in each compartment in an ODE. Then we define $\mathcal{F}_{i}(x)$ to be the rate at which tokens move into infected places. For the variable arc weight PN, this rate refers to the arc weights. This is represented in matrix form by
$\mathcal{F}_{1,...,T}(x)$ $=$ $\begin{bmatrix}w(t_{N_{j,...,k}},I_{1})&w(t_{N_{j,...,k}},I_{1})&\cdots&w(t_{N_{j,...,k}},I_{1})\\ w(t_{N_{j,...,k}},I_{2})&w(t_{N_{j,...,k}},I_{2})&\cdots&w(t_{N_{j,...,k}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ w(t_{N_{j,...,k}},I_{T})&w(t_{N_{j,...,k}},I_{T})&\cdots&w(t_{N_{j,...,k}},I_{T}))\end{bmatrix}.$ (2)
Looking at Figure 1 we find that the elements utilized in $\mathcal{F}$ are outlined in red. This is continued in all the other Petri net models in this paper to make it easier to find what elements fit where by inspection. For $\mathcal{F}_{1}$ the element $w(t_{N_{j,...,k}},I_{1})=\beta_{1}$. For $\mathcal{F}_{2}$ the element $w(t_{N_{j,...,k}},I_{2})=\beta_{2}$. For $\mathcal{F}_{3}$ no transmission occurs directly into place $I_{3}$, so the element $w(t_{N_{j,...,k}},I_{3})=0$. Thus
$\mathcal{F}_{1,2,3}(x)$ $=$ $\begin{bmatrix}\beta_{1}&\beta_{1}&\beta_{1}\\ \beta_{2}&\beta_{2}&\beta_{2}\\ 0&0&0\end{bmatrix}.$
We define $w(I_{i},t_{j,...,k})$ as the summation of arc weights of output arcs from infected place $I_{i}$ to transitions $t_{j,...,k}$ for VAPNs and the summation of rate equations of all arcs between infected place $I_{i}$ to transitions $t_{j,...,k}$. In other words, this represents adding up all arcs leaving infected place $I_{i}$. We define $w(t_{I_{i}},I_{j})$ as the arc weights of arcs from transition to infected places where the arcs leading into the transition come from infected place $I_{i}$, and this arc weight is equal to the arc weight of the arc from $t_{I_{i}}$ to $I_{j}$. Lastly, we define $w(t_{N_{m},...,N_{l}},I_{j})$ as the sum of the arc weights of arcs from transitions to infected place $I_{j}$ where the transitions $t_{N_{m},...,N_{l}}$ have no arcs leading into the transitions. We define $V^{-}_{i}(x)$ to be the arc weight of tokens leaving infected place $i$ and $V^{+}_{i}(x)$ to be the arc weight of tokens going into infected place $i$ outside of coming from non-infected places. In terms of what the terms $\mathcal{V}^{-}_{i}(x)$ and $\mathcal{V}^{+}_{i}(x)$ represent, it is the rate of people leaving the infected place $i$ and the rate of people going into infected place $i$ from anything other than coming from non-infected places, respectively. These terms combine to yield, $\mathcal{V}_{i}(x)=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x)$. For Petri nets, this can be expressed through the matrix
$\mathcal{V}_{1,...,T}(x)=\begin{bmatrix}w(I_{1},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{1})&-w(t_{I_{2}},I_{1})&\cdots&-w(t_{I_{T}},I_{1})\\ -w(t_{I_{1}},I_{2})&w(I_{2},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{2})&\cdots&-w(t_{I_{T}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ -w(t_{I_{1}},I_{T})&-w(t_{I_{2}},I_{T})&\cdots&w(I_{T},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{T})\end{bmatrix}.$ (3)
Looking at Figure 1 we find that the elements utilized in $\mathcal{V}$ are outlined in green. This is continued in all the other Petri net models in this paper to make it easier to find what elements fit in $\mathcal{V}$ by inspection. For $\mathcal{V}_{1}$ the elements $w(I_{1},t_{j,...,k})=\gamma_{1}$, $w(t_{N_{m},...,N_{l}},I_{1})=0$, $-w(t_{I_{2}},I_{1})=0$, and $-w(t_{I_{3}},I_{1})=0$. For $\mathcal{V}_{2}$ the elements $-w(t_{I_{1}},I_{2})=0$, $w(I_{2},t_{j,...,k})=\gamma_{2}$, $w(t_{N_{m},...,N_{l}},I_{2})=0$, and $--w(t_{I_{3}},I_{2})=0$. For $\mathcal{V}_{3}$ the elements $-w(t_{I_{1}},I_{T})=-\delta_{1}$, $-w(t_{I_{2}},I_{3})=-\delta_{2}$, $w(I_{3},t_{j,...,k})=\varepsilon$, and $w(t_{N_{m},...,N_{l}},I_{3})=0$. Thus,
$\mathcal{V}_{1,2,3}(x)$ $=$ $\begin{bmatrix}\gamma_{1}&0&0\\ 0&\gamma_{2}&0\\ -\delta_{1}&-\delta_{2}&\varepsilon\end{bmatrix}.$
The disease free equilibrium (DFE) for a system with $L$ places is defined as ($P_{1}^{*},....,P_{L}^{*}$), the state at which the system remains absent of the disease. For the Petri nets the DFE can be found by first setting the token level of all infected places to zero. Then, the arc weights of arcs into and out of a given non-infected place $P_{i}$ are summed, with the the arc weights of arcs going out of $P_{i}$ being multiplied by -1. The resulting expression from these steps is then set to zero. From here, a system of equations is formed that can be solved for $P_{1}^{*},....,P_{L}^{*}$. Additionally, the total population, usually denoted as $N$, is defined as $\mathcal{N}+\varepsilon$, where $\mathcal{N}$ is the summation of tokens in all places, and $\varepsilon$ is the error from rounding in a discrete token level Petri net. In a continuous token level Petri net $\varepsilon=0$. The difference in total population between $N$ and $\mathcal{N}+\varepsilon$ comes into play when doing actual simulations of models not when finding $R_{0}$ using NGMPN method. From here, a system of equations is formed where the DFE token levels labeled for each non-infected place can be solved for.
Once these $\mathcal{F}_{i}(x)$, $\mathcal{V}_{i}(x)$, and DFE values are obtained, they are then converted into the respective $F$ (the transmission matrix) and $V$ (the transition matrix) Jacobian matrix by finding the rate of change with respect to $I_{1}$ for column one, $I_{2}$ for column two, $\cdots$, $I_{T}$ for column $T$, at the DFE, for the respective arc weights defined as functions. Note that Petri nets can be in discrete or continuous time. Thus, the rate of change would be the difference quotient or derivative, respectively.
Then, using the Petri net based $F$ and $V$ values, the same Next-Generation Matrix steps of finding the inverse of $V$ and finding the dominant eigenvalues of $FV^{-1}$ can be done to find the Petri net $R_{0}$. If the values of $F,V$ and the DFE of two systems are equal, then the resulting $R_{0}$ values are equal. However, the way that parameters interact within an ODE model and a PN model are different. We will show the process in detail with various examples with different functionality.
Addressing the assumptions (A1)-(A5) made in van den Driessche and Watmough [6], we interpret them for Petri nets.
  1. (A1)
    If a place $P_{i}$ contains a non-negative number of tokens $t_{i}\geq 0$, then the number of tokens that can be added to or removed from any place by any transition is non-negative. Specifically, if a transition $t_{j}$ is enabled (based on its input places and corresponding markings), the number of tokens removed from the input places and added to the output places must be non-negative. By the definition of Petri nets, the assumption is inherently preserved across all Petri net structures. This is due to the stipulation that the number of tokens in each place must remain non-negative at all times. Additionally, transitions in a Petri net operate in a manner that ensures the non-negativity of tokens, effectively maintaining this essential characteristic.
  2. (A2)
    If the marking of a place $P_{i}$ is zero (i.e., $t_{i}=0$), then no transition can consume tokens from place $P_{i}$. Specifically, if a place is empty (i.e., no tokens are present in the corresponding place), no transition can have it as an input. This ensures that no ”removal” of individuals (tokens) from an empty place occurs. Again the definition of Petri net preserves this assumption for all Petri nets as a transition can not fire if an input place has less tokens than the corresponding arc weight. If a place has no tokens (individuals), no transition can remove tokens from that place.
  3. (A3)
    If place $P_{i}$ corresponds to a non-infectious place, then no transition can place tokens into this place from a transition that represents infection occurring or remaining. No transition should be able to fire that would place tokens into non-infectious places from a transition that represents infection occurring or remaining.
  4. (A4)
    If the marking of a place $P_{i}$ indicates a disease-free population (i.e., there are no tokens representing infected individuals in this marking), then no transition can fire that would result in tokens (representing infected individuals) being moved to an infected place. If a place is has a disease-free marking, no transitions can place tokens in it that correspond to infected individuals (i.e., no ”infection” transition should have this place as an output if it is currently disease-free).
  5. (A5)
    If $\mathcal{F}(x)$ is set to 0, then all eigenvalues of $Df(x)$ have negative real parts. In terms of Petri nets, this means that if there are no transitions that change the marking of infected places coming from non-infected places, then the related infected place’s markings will go to zero over time. If a few infected tokens are introduced into the system, the infected place’s markings will still go to zero over time. No new infections means only infections leaving, thus the resulting $Df(x)<0$.
4 Case Studies
We have selected a range of ODE and corresponding PN example models to show how the Next-Generation Matrix for Petri nets (NGMPN) approach to finding $R_{0}$. Each model either adds additional compartments/places or new dynamics to provide a range of examples for people to draw on if they wish to implement the NGMPN on a new model.
4.1 SIRS
We start with a simple example of an $SIRS$ system, a small variation on the Kermack-McKendrick model [19]. This shows a straightforward application of the NGMPN process for a foundational epidemiological model. The compartments $S,I,R$ represent susceptible, infected, and recovered, respectively. The parameters of this system $\beta,\delta,$ and $\gamma$ represent the rate of re-susceptibility, rate of infection, and rate of recovery, respectively.
$\frac{dS}{dt}$ $=$ $-\beta SI+\delta R$ (4)
$\frac{dI}{dt}$ $=$ $\beta SI-\gamma I$ (5)
$\frac{dR}{dt}$ $=$ $\gamma I-\delta R$ (6)
The ODE system can be mapped into two Petri net models, one with variable arc weights and minimal arcs yielding Figure 2(a), and another using the methods laid out by Baez and Biamonte [18], yielding Figure2(b).
Figure 2: Two Petri net implementations of the SIRS ODE model given in Equations 4-6. Option (a) uses a discrete-time, variable arc weight PN. Option (b) uses a continuous-time, stochastic Markov Chain PN. The arcs and arc weights used in forming matrix $F_{i}(x)$ are outlined and highlighted in red, and the arcs and arc weights used in forming matrix $V_{i}(x)$ are outlined and highlighted in green.
The first step in finding the $R_{0}$ of the PN systems in Figure 2 is to obtain the Disease Free Equilibrium (DFE) by setting the token levels to zero in the infected places, setting the net arc weight of non-infected places to 0, and solving for the token values in non-infected places $(S^{*},I^{*},R^{*})=(S^{*},0,0)$. Then the transmission matrix starts with finding $\mathcal{F}_{i}(x)$, which can be found by looking at the arc weights of arcs going into the infected compartments as shown in Equation 2. These arcs are outlined in red in Figure 2. Note that $i$ is with respect to the corresponding infected compartment, which we number. For this example, the $I$ compartment is 1. From here we find $w(t_{N_{j,...,k}},I_{1})=\beta SI$ for the VAPN in Figure 2(a) and $w(t_{N_{j,...,k}},I_{1})=2\beta SI-\beta SI$ for the SPN in Figure 2(b). Thus,
$\mathcal{F}_{1}(x)$ $=$ $\begin{bmatrix}\beta SI\end{bmatrix}$
for the Petri net layout as depicted in Figure 2(a), and
$\mathcal{F}_{1}(x)$ $=$ $\begin{bmatrix}2\beta SI-\beta SI\end{bmatrix}.$
for the Petri net layout as depicted in Figure 2(b). From this, we find the rate of change with respect to the infected place $I$, and we evaluate at the DFE to have
$F$ $=$ $\begin{bmatrix}\beta\end{bmatrix}.$
for both layouts. For the transition matrix, we start with $\mathcal{V}$, and we look at arc weights of arcs leaving the infected places as shown in Equation 3. These arcs are outlined in green in Figure 2. The terms $w(I_{1},t_{j,...,k})=\gamma I$ and $w(t_{N_{m},...,N_{l}}=0$ for both the VAPN in Figure 2(a) and the SPN in Figure 2(b). Thus yielding for both layouts
$\mathcal{V}_{1}(x)$ $=$ $\begin{bmatrix}\gamma I\end{bmatrix}.$
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this we find the inverse:
$V$ $=$ $\begin{bmatrix}\gamma\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\gamma}\end{bmatrix}.$
Now that we have the $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\frac{\beta}{\gamma}$
This $R_{0}$ value of both layouts of $\frac{\beta}{\gamma}$ is analytically equivalent to the ODE $R_{0}$ value.
4.2 SEIR
This second example introduces an exposed compartment for diseases with delayed onset, a version of the model was laid out by Cooke in 1966 [20]. The additional compartment $E$ represents exposed, usually meaning a person is infected with a disease but can’t infect others yet. The parameters $\Pi,\beta,\mu,\eta,$ and $\alpha$ represent the birth/immigration rate, the infection rate, the death/emigration rate, the onset of infectiousness/symptoms rate, and the recovery rate, respectively.
$\frac{dS}{dt}$ $=$ $\Pi-\beta SI-\mu S$ (8)
$\frac{dE}{dt}$ $=$ $\beta SI-\eta E-\mu E$ (9)
$\frac{dI}{dt}$ $=$ $\eta E-\alpha I-\mu I$ (10)
$\frac{dR}{dt}$ $=$ $\alpha I-\mu R$ (11)
The Equations 8-11 are mapped to a VAPN (Figure 3(a)) and stochastic Petri net (Figure 3(b)).
Figure 3: Two Petri net implementations of the SEIR ODE model given in Equations 8-11. Option (a) uses a discrete-time, variable arc weight PN. Option (b) uses a continuous-time, stochastic Markov Chain PN. The arcs and arc weights used in forming matrix $F_{i}(x)$ are outlined and highlighted in red, and the arcs and arc weights used in forming matrix $V_{i}(x)$ are outlined and highlighted in green.
The first step in finding the $R_{0}$ of the PN systems in Figure 3 is to obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E^{*},I^{*},R^{*})=(\frac{\Pi}{\mu},0,0,0,0)$. Then the matrix $F_{i}(x)$ can be found by looking at the arc weights of arcs going into the infected compartments, outlined in red in Figure 3 and as shown in Equation 2. Note that $i$ is with respect to the corresponding infected compartment and the corresponding row in the matrix, which we number. For this example, the $E$ compartment is 1 and the $I$ compartment is 2. The terms for the VAPN in Figure 3(a) within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,...,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,...,k}},I_{1})=0$. Similarly the terms for the SPN in Figure 3(b) within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,...,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,...,k}},I_{1})=\beta SI-\beta SI=0$. Thus,
$\mathcal{F}_{1,2}(x)$ $=$ $\begin{bmatrix}\beta SI&\beta SI\\ 0&0\end{bmatrix},$
for VAPN as shown in Figure 3(a), and
$\mathcal{F}_{1,2}(x)$ $=$ $\begin{bmatrix}\beta SI&\beta SI\\ \beta SI-\beta SI&\beta SI-\beta SI\end{bmatrix},$
for the stochastic PN as shown in Figure 3(b).
From this we find the rate of change with respect to $E$ in the first column and $I$ in the second column, and evaluate at the DFE. This gives the same result for both Petri net options:
$F$ $=$ $\begin{bmatrix}0&\beta\frac{\Pi}{\mu}\\ 0&0\\ \end{bmatrix}.$
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another, as outlined in green in Figure 3. Then for $\mathcal{V}_{1}$ the terms $w(I_{1},t_{j,...,k})=\eta E+\mu E$, $w(t_{N_{m},...,N_{l}}=0,I_{1})$, and $-w(t_{I_{2}},I_{1})=0$. For $\mathcal{V}_{2}$ the terms $-w(t_{I_{1}},I_{2})=-\eta E$, $w(I_{2},t_{j,...,k})=\alpha I+\mu I$, and $w(t_{N_{m},...,N_{l}},I_{2})=0$. Both $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are the same for the VAPN and SPN in Figures 3(a) and 3(b) respectively. Thus,
$\mathcal{V}_{1,2}(x)$ $=$ $\begin{bmatrix}\eta E+\mu E&0\\ -\eta E&\alpha I+\mu I\end{bmatrix}.$
We then find the rate of change with respect to the infected place $E$ in the first column and $I$ in the second column, and we evaluate it at the DFE. From this we find the inverse to give
$V$ $=$ $\begin{bmatrix}\eta+\mu&0\\ -\eta&\alpha+\mu\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\eta+\mu}&0\\ \frac{a}{(n+u)(a+u)}&\frac{1}{a+\mu}\end{bmatrix}.$
Now that we have the $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\frac{\beta\Pi\eta}{\mu(\alpha+\mu)(\eta+\mu)}$
This is analytically the same $R_{0}$ expression that can be derived for the ODE using the next-generation matrix method.
4.3 SEEIR
This third example introduces parallel paths of infected compartments. The model and subsequent $R_{0}$ value was outlined by Diekmann [5]. The parameters $\mu,\beta,p,v_{1},v_{2},$ and $\gamma$ represent birth/death rate, transmission rate, rate of exposure to infection, rate of becoming infectious for $E_{1}$, rate of becoming infectious for $E_{2}$, and rate of recovery, respectively.
$\frac{dS}{dt}$ $=$ $\mu N-\beta\frac{SI}{N}-\mu S$ (13)
$\frac{dE_{1}}{dt}$ $=$ $p\beta\frac{SI}{N}-v_{1}E_{1}-\mu E_{1}$ (14)
$\frac{dE_{2}}{dt}$ $=$ $(1-p)\beta\frac{SI}{N}-v_{2}E_{2}-\mu E_{2}$ (15)
$\frac{dI}{dt}$ $=$ $v_{1}E_{1}+v_{2}E_{2}-\gamma I-\mu I$ (16)
$\frac{dR}{dt}$ $=$ $\gamma I-\mu R$ (17)
The Equations 13-17 are mapped to a variable arc weight Petri net to yield Figure 4. We will analyze only the Variable Arc Weight Petri net (VAPN) implementation going forward for brevity and since it is the more novel approach.
Figure 4: This SEEIR Petri net model is mapped directly from Equations 13-17. The arc weights used in forming matrix $F_{i}(x)$ are outlined in red and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
The first step in finding the $R_{0}$ of the PN system in Figure 4 is to obtain the DFE by setting the token levels of the infected places to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E_{1}^{*},E_{2}^{*},I^{*},R^{*})=(S^{*},0,0,0,0)$. Note since $N=S+E_{1}+E_{2}+I+R$, $S^{*}=N$, where $N$ is the total population. Then the matrix $F_{i}(x)$ can be found by looking at the arc weights of arcs going into the infected compartments of $E_{1},E_{2},$ and $I$ as shown in Equation 2. Note that $i$ refers to the corresponding infected compartment and the corresponding row in the matrix, which we have numbered. For this example, the $E_{1}$ compartment is 1, the $E_{2}$ compartment is 2 and the $I$ compartment is 3.
$\mathcal{F}_{1,2,3}(x)$ $=$ $\begin{bmatrix}p\beta\frac{SI}{N}&p\beta\frac{SI}{N}&p\beta\frac{SI}{N}\\ (1-p)\beta\frac{SI}{N}&(1-p)\beta\frac{SI}{N}&(1-p)\beta\frac{SI}{N}\\ 0&0&0\end{bmatrix}.$
From this, we find the rates of change with respect to $E_{1}$ in the first column, $E_{2}$ in the second column, $I$ in the third column, and we evaluate them at the DFE to give
$F$ $=$ $\begin{bmatrix}0&0&p\beta\\ 0&0&(1-p)\beta\\ 0&0&0\end{bmatrix}.$
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another as shown in Equation 3. This yields
$\mathcal{V}_{1,2,3}(x)$ $=$ $\begin{bmatrix}v_{1}E_{1}+\mu E_{1}&0&0\\ 0&v_{2}E_{2}+\mu E_{2}&0\\ -v_{1}E_{1}&-v_{2}E_{2}&\gamma I+\mu I\end{bmatrix}.$
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this we find the inverse to give
$V$ $=$ $\begin{bmatrix}v_{1}+\mu&0&0\\ 0&v_{2}+\mu&0\\ -v_{1}&-v_{2}&\gamma+\mu\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{v_{1}+\mu}&0&0\\ 0&\frac{1}{v_{2}+\mu}&0\\ \frac{v_{1}}{(v_{1}+\mu)(\gamma+\mu)}&\frac{v_{2}}{(v_{2}+\mu)(\gamma+\mu)}&\frac{1}{\gamma+\mu}\end{bmatrix}.$
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\left(\frac{pv_{1}}{v_{1}+\mu}+\frac{(1-p)v_{2}}{v_{2}+\mu}\right)\frac{\beta}{\gamma+\mu}$
This is analytically the same $R_{0}$ expression found in Equation 2.10 of [5].
4.4 Basic COVID Model
This next example was laid out as a basic COVID-19 model to help with forecasting once the appropriate compartments for the disease had been classified [21]. This example introduces more complex equations for $F$, more interactions involved in $V$, and application to a specific disease. By applying our method to a compartmental COVID-19 model, we demonstrate its practical utility in analyzing contemporary and complex epidemiological challenges, where multiple infectious states with varying transmission dynamics play a crucial role. The compartments $S,E,I_{a},I_{s},I_{h},$ and $R$ represent susceptible, exposed, infected asymptomatic, infected symptomatic, infected hospitalized, and recovered, respectively. The parameters $\beta_{a},\beta_{S},$ and $\beta_{h}$ represent the rate of infection due to asymptomatic, symptomatic, and hospitalized individuals, respectively. The parameters $\sigma,r,\gamma_{a},\gamma_{s},\varphi_{S},\delta_{s},\gamma_{h},$ and $\delta_{h}$ represent rate of becoming infectious, the percentage of people who are asymptomatic, rate of recovery for asymptomatic infected, rate of recovery for symptomatic infected, rate of hospitalization of symptomatic, death rate of symptomatic, rate of recovery for hospitalized and death rate of hospitalized, respectively.
$\frac{dS}{dt}$ $=$ $-\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S$ (18)
$\frac{dE}{dt}$ $=$ $\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S-\sigma E$ (19)
$\frac{dI_{a}}{dt}$ $=$ $r\sigma E-\gamma_{a}I_{a}$ (20)
$\frac{dI_{S}}{dt}$ $=$ $(1-r)\sigma E-(\varphi_{s}+\gamma_{s}+\delta_{s})I_{s}$ (21)
$\frac{dI_{h}}{dt}$ $=$ $\varphi_{s}I_{s}-(\gamma_{h}+\delta_{h})I_{h}$ (22)
$\frac{dR}{dt}$ $=$ $\gamma_{a}I_{a}+\gamma_{s}I_{s}+\gamma_{h}I_{h}$ (23)
The Equations 18-23 are mapped to a variable arc weight Petri net to yield Figure 5.
Figure 5: A variable arc weight COVID Petri net mapped from Equations 18-23. The arc weights used in forming matrix $F_{i}(x)$ are outlined in red and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
From Figure 5, we can see that there are four infected places ($E,I_{a},I_{S},I_{h}$) which are the first, second, third, and fourth infected compartments, respectively. Our resulting $F$ and $V$ matrices will therefore be four by four. The disease free equilibrium is given by $(S^{*},E^{*},I_{a}^{*},I_{s}^{*},I_{h}^{*},R^{*})=(S^{*},0,0,0,R^{*})$. We defined $N=S+E+I_{a}+I_{S}+I_{h}+R$, so $N^{*}=S^{*}+R^{*}$. By inspecting the PN model we see that the term $\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S$ is the only way the Susceptible place leads to an infected place. The transmission matrix starts with the resulting $\mathcal{F}$ using Equation 2 is
$\mathcal{F}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$
Taking the partial derivative with respect $E,I_{a},I_{S},I_{h}$ for the first, second, third, and fourth columns, respectively, gives us
$F$ $=$ $\begin{bmatrix}0&\frac{\beta_{a}S^{*}}{N^{*}}&\frac{\beta_{S}S^{*}}{N^{*}}&\frac{\beta_{h}S^{*}}{N^{*}}\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$
where $S^{*},N^{*}$ are the $S,N$ values at the DFE.
For $V$, the transition matrix, we start by finding $\mathcal{V}$ based on the arc weights between infected places and the arc weights leaving the infected places as shown in Equation 3. Thus
$\mathcal{V}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\sigma E&0&0&0\\ -r\sigma E&\gamma_{a}I_{a}&0&0\\ -(1-r)\sigma E&0&\varphi_{S}I_{S}+\gamma_{S}I_{S}+\delta_{S}I_{S}&0\\ 0&0&-\varphi_{S}I_{S}&\gamma_{h}I_{H}\end{bmatrix}$
Then taking the partial derivative with respect $E,I_{a},I_{S},I_{h}$ for the first, second, third, and fourth columns, respectively, gives us
$V$ $=$ $\begin{bmatrix}\sigma&0&0&0\\ -r\sigma&\gamma_{a}&0&0\\ -(1-r)\sigma&0&\varphi_{S}+\gamma_{S}+\delta_{S}&0\\ 0&0&\varphi_{S}&\gamma_{h}\end{bmatrix}$
Hence,
$R_{0}=\varrho(FV^{-1})=\left(\left(\frac{\beta_{a}r}{\gamma_{a}}\right)+\left(\frac{\beta_{s}(1-r)}{\varphi_{s}+\gamma_{s}+\delta_{s}}\right)+\left(\frac{\beta_{h}(1-r)\varphi_{s}}{(\varphi_{s}+\gamma_{s}+\delta_{s})(\gamma_{h}+\delta_{h})}\right)\right)\left(\frac{S^{*}}{N^{*}}\right).$
The values for the PN $F$ and $V$ are equivalent to that values for the ODE $F$ and $V$ and the resulting $R_{0}$ matches the $R_{0}$ for the ODE system in the Gumel paper [21] seen in Equations (2) and (3).
4.5 Nonlinear System
The paper by Rohith et al. [22] involves nonlinear compartment-to-compartment dynamics. These types of dynamics are more applicable when wanting to involve the dynamics of social aspects of a disease spread, drug pharmacokinetics, or disease cell spread within an individual. The analysis of the nonlinear system underscores the robustness of the NGMPN framework, demonstrating its applicability even when departing from standard mass-action assumptions to incorporate more realistic, behavior-driven transmission dynamics. The parameters $\mu,\beta,\alpha,\sigma,$ and $\gamma$ represent the birth/death rate, the per capita contact rate, the psychological or inhibitory effect, rate of infectiousness, and the recovery rate, respectively.
$\frac{dS}{dt}$ $=$ $\mu-\frac{\beta SI}{1+\alpha I^{2}}-\mu S$ (24)
$\frac{dE}{dt}$ $=$ $\frac{\beta SI}{1+\alpha I^{2}}-\sigma E-\mu E$ (25)
$\frac{dI}{dt}$ $=$ $\sigma E-\gamma I-\mu I$ (26)
$\frac{dR}{dt}$ $=$ $\gamma I-\mu R$ (27)
Figure 6: Non-linear SEIR Variable Arc Weight PN Model
From Figure 6 we can see that there are two infected places $(E,I)$. Thus our resulting $F$ and $V$ matrices will be two by two. Then, from inspecting the PN model, we see that $\frac{\beta SI}{1+\alpha I^{2}}$ is the only way that the new infected population enters. Thus the resulting $\mathcal{F}$ is
$\mathcal{F}_{1,2}(x)$ $=$ $\begin{bmatrix}\frac{\beta SI}{1+\alpha I^{2}}&\frac{\beta SI}{1+\alpha I^{2}}\\ 0&0\end{bmatrix}.$
Then finding the DFE of the PN system to be $(S^{*},E^{*},I^{*},R^{*})=(1,0,0,0)$. We then take the partial derivative of each element of $\mathcal{F}_{1,2}$ with respect to $E,I$ for the first and second columns, respectively, to give the transmission matrix $F$,
$F$ $=$ $\begin{bmatrix}0&\beta\\ 0&0\end{bmatrix}.$
For the transition matrix $V$, we start by finding $\mathcal{V}$ based on the arc weights between infected places and the arc weights leaving the infected places yielding
$\mathcal{V}_{1,2}(x)$ $=$ $\begin{bmatrix}\sigma E+\mu E&0\\ -\sigma E&\gamma I+\mu I\end{bmatrix}.$
Then, taking the partial derivative with respect to $E,I$ for the first and second column gives us
$V$ $=$ $\begin{bmatrix}\sigma+\mu&0\\ -\sigma&\gamma+\mu\end{bmatrix}.$
Then we find the inverse of $V$,
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\sigma+\mu}&0\\ \frac{\sigma}{(\sigma+\mu)(\gamma+\mu)}&\frac{1}{\gamma+\mu}\end{bmatrix}.$
Then with the values of $F$ and $V^{-1}$ we can find the dominant eigen value of $(FV^{-1})$:
$R_{0}=\varrho(FV^{-1})=\frac{\sigma\beta}{(\sigma+\mu)(\gamma+\mu)},$
which is the same $R_{0}$ value found in the Rohith paper [22].
4.6 Patch System Overview
The paper Multi-patch and multi-group epidemic models: a new framework by Bichara and Iggidr [23] presents a multi-patch and multi-group model for infectious disease transmission. The model accommodates interactions occurring across an arbitrary number of patches and groups where the infection propagates. The compartments $S_{i},E_{i},I_{i},R_{i}$ still represent susceptible, exposed, infected, and recovered, respectively, just now for patch $i$. Similarly, the parameters $\Pi,\beta,\mu,\eta,\upsilon,\gamma,$ and $\delta$ represent birth/immigration, rate of infection, natural death rate, loss of immunity rate, rate of becoming infectious, rate of recovery and rate of disease related death, respectively. The parameter $\beta$ is a patch specific term while $\Pi,\mu,\eta,\upsilon,\gamma,$ and $\delta$ are group specific terms. People of group $i$ spend on average some time in patch $j$,$j=1,...,v$. The susceptible, exposed, infected, and recovered populations of group $i$ spend $m_{ij},n_{ij},p_{ij},$ and $q_{ij}$ proportion of time in patch $j$ for $j=1,...,v$. , , and proportion of times respectively in Patch j, for . The ODEs are outlined here
$\frac{dS_{i}}{dt}$ $=$ $\Pi_{i}-\sum_{j=1}^{v}\beta_{j}m_{ij}S_{i}\frac{\sum_{k=1}^{u}p_{kj}I_{k}}{\sum_{k=1}^{u}(m_{kj}S_{k}+n_{kj}E_{k}+p_{kj}I_{k}+q_{kj}R_{k})}-\mu_{i}S_{i}+\eta_{i}R_{i}$ (28)
$\frac{dE_{i}}{dt}$ $=$ $\sum_{j=1}^{v}\beta_{j}m_{ij}S_{i}\frac{\sum_{k=1}^{u}p_{kj}I_{k}}{\sum_{k=1}^{u}(m_{kj}S_{k}+n_{kj}E_{k}+p_{kj}I_{k}+q_{kj}R_{k})}-\upsilon_{i}E_{I}-\mu_{i}E_{i}$ (29)
$\frac{dI_{i}}{dt}$ $=$ $\upsilon_{i}E_{I}-\gamma_{i}I_{i}-\delta_{i}I_{i}-\mu_{i}I_{i}$ (30)
$\frac{dR_{i}}{dt}$ $=$ $\gamma_{i}I_{i}-\eta_{i}R_{i}-\mu_{i}R_{i}.$ (31)
4.6.1 Two Patch System
We take the ODEs from Bichara and Iggidr [23] Equations 28-31 and lay them out explicitly for two patches and two groups. Though the ODE and corresponding PN model only are for this small patch and group size, this model could be extended to any size with the interactions between any two given patches being equivalent to the system laid out below. The multi-patch model highlights the scalability of our approach to spatially structured populations, providing a tool to investigate the impact of geographic factors and population mobility on disease spread.
$\frac{dS_{1}}{dt}$ $=$ $\Pi_{1}-\left(\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (32)
$\beta_{2}m_{12}S_{1}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\mu_{1}S_{1}+\eta_{1}R_{1}$
$\frac{dS_{2}}{dt}$ $=$ $\Pi_{2}-\left(\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (33)
${}_{22}\beta_{2}m_{22}S_{2}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\mu_{2}S_{2}+\eta_{2}R_{2}$
$\frac{dE_{1}}{dt}$ $=$ $\left(\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (34)
$\beta_{2}m_{12}S_{1}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{k}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\upsilon_{1}E_{1}-\mu_{1}E_{1}$
$\frac{dE_{2}}{dt}$ $=$ $\left(\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}+\right....$ (35)
$\beta_{2}m_{22}S_{2}\left.\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{k}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}\right)...$
$-\upsilon_{2}E_{2}-\mu_{2}E_{2}$
$\frac{dI_{1}}{dt}$ $=$ $\upsilon_{1}E_{1}-\gamma_{1}I_{1}-\delta_{1}I_{1}-\mu_{1}I_{1}$ (36)
$\frac{dI_{2}}{dt}$ $=$ $\upsilon_{2}E_{2}-\gamma_{2}I_{2}-\delta_{2}I_{2}-\mu_{2}I_{2}$ (37)
$\frac{dR_{1}}{dt}$ $=$ $\gamma_{1}I_{1}-\eta_{1}R_{1}-\mu_{1}R_{1}$ (38)
$\frac{dR_{2}}{dt}$ $=$ $\gamma_{2}I_{2}-\eta_{2}R_{2}-\mu_{i}R_{2}$ (39)
Figure 7: Patch Variable Arc Weight PN Model
To simplify, we define $\Theta_{1}=\beta_{1}m_{11}S_{1}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}$,
$\Theta_{2}=\beta_{2}m_{12}S_{1}\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}$,
$\Theta_{3}=\beta_{1}m_{21}S_{2}\frac{p_{11}I_{1}+p_{21}I_{2}}{(m_{11}S_{1}+n_{11}E_{1}+p_{11}I_{1}+q_{11}R_{1})+(m_{21}S_{2}+n_{21}E_{2}+p_{21}I_{2}+q_{21}R_{2})}$,
$\Theta_{4}=\beta_{2}m_{22}S_{2}\frac{p_{12}I_{1}+p_{22}I_{2}}{(m_{12}S_{1}+n_{12}E_{1}+p_{12}I_{1}+q_{12}R_{1})+(m_{22}S_{2}+n_{22}E_{2}+p_{22}I_{2}+q_{22}R_{2})}$.
Then we obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the non-infected places token values $(S_{1}^{*},S_{2}^{*},E_{1}^{*},E_{2}^{*},I_{1}^{*},I_{2}^{*},R_{1}^{*},R_{2}^{*})=(\frac{\Pi_{1}}{\mu_{1}},\frac{\Pi_{2}}{\mu_{2}},0,0,0,0,0,0)$. Then the matrix $\mathcal{F}$ by looking at the arc weights of arcs going into the infected compartments giving us
$\mathcal{F}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\Theta_{1}+\Theta_{2}&\Theta_{1}+\Theta_{2}&\Theta_{1}+\Theta_{2}&\Theta_{1}+\Theta_{2}\\ \Theta_{3}+\Theta_{4}&\Theta_{3}+\Theta_{4}&\Theta_{3}+\Theta_{4}&\Theta_{3}+\Theta_{4}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}.$
From this we find the rate of change with respect to $E_{1}$ in the first column, $E_{2}$ in the second column, $I$ in the third column, and evaluate at the DFE to give
$F$ $=$ $\begin{bmatrix}0&0&\frac{\beta_{1}m_{11}p_{11}S_{1}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{12}p_{12}S_{1}}{m_{12}S_{1}+m_{22}S_{2}}&\frac{\beta_{1}m_{11}p_{12}S_{1}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{12}p_{22}S_{1}}{m_{12}S_{1}+m_{22}S_{2}}\\ 0&0&\frac{\beta_{1}m_{21}p_{11}S_{2}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{22}p_{12}S_{2}}{m_{12}S_{1}+m_{22}S_{2}}&\frac{\beta_{1}m_{21}p_{21}S_{2}}{m_{11}S_{1}+m_{21}S_{2}}+\frac{\beta_{2}m_{22}p_{22}S_{2}}{m_{12}S_{1}+m_{22}S_{2}}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}.$
For $\mathcal{V}$, we look at arc weights of arcs leaving the infected places and transitioning from one infected place to another yielding
$\mathcal{V}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\nu_{1}E_{1}+\mu_{1}E_{1}&0&0&0\\ 0&\nu_{2}E_{2}+\mu_{2}E_{2}&0&0\\ -\nu_{1}E_{1}&0&\gamma_{1}I_{1}+\delta_{1}I_{1}+\mu_{1}I_{1}&0\\ 0&-\nu_{2}E_{2}&0&\gamma_{2}I_{2}+\delta_{2}I_{2}+\mu_{2}I_{2}\end{bmatrix}.$
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this we find the inverse to give
$V$ $=$ $\begin{bmatrix}\nu_{1}+\mu_{1}&0&0&0\\ 0&\nu_{2}+\mu_{2}&0&0\\ -\nu_{1}&0&\gamma_{1}+\delta_{1}+\mu_{1}&0\\ 0&-\nu_{2}&0&\gamma_{2}+\delta_{2}+\mu_{2}\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\nu_{1}+\mu_{1}}&0&0&0\\ 0&\frac{1}{\nu_{2}+\mu_{2}}&0&0\\ \frac{\nu_{1}}{(\nu_{1}+\mu_{1})(\gamma_{1}+\delta_{1}+\mu_{1})}&0&\frac{1}{\gamma_{1}+\delta_{1}+\mu_{1}}&0\\ 0&\frac{\nu_{2}}{(\nu_{2}+\mu_{2})(\gamma_{2}+\delta_{2}+\mu_{2})}&0&\frac{1}{\gamma_{2}+\delta_{2}+\mu_{2}}\end{bmatrix}.$
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and determining the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}F\cdot V^{-1}\end{bmatrix}=$
$\varphi{\tiny\begin{bmatrix}\frac{\left(\frac{\beta_{1}m_{{11}}p_{{11}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{12}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{1}}{\left(v_{1}+\mu_{1}\right)\left(\gamma_{1}+\delta_{1}+\mu_{1}\right)}&\frac{\left(\frac{\beta_{1}m_{{11}}p_{{12}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{22}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{2}}{\left(v_{2}+\mu_{2}\right)\left(\gamma_{2}+\delta_{2}+\mu_{2}\right)}&\frac{\frac{\beta_{1}m_{{11}}p_{{11}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{12}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{1}+\delta_{1}+\mu_{1}}&\frac{\frac{\beta_{1}m_{{11}}p_{{12}}S_{1}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{12}}p_{{22}}S_{1}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{2}+\delta_{2}+\mu_{2}}\\ \frac{\left(\frac{\beta_{1}m_{{21}}p_{{11}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{12}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{1}}{\left(v_{1}+\mu_{1}\right)\left(\gamma_{1}+\delta_{1}+\mu_{1}\right)}&\frac{\left(\frac{\beta_{1}m_{{21}}p_{{21}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{22}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}\right)v_{2}}{\left(v_{2}+\mu_{2}\right)\left(\gamma_{2}+\delta_{2}+\mu_{2}\right)}&\frac{\frac{\beta_{1}m_{{21}}p_{{11}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{12}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{1}+\delta_{1}+\mu_{1}}&\frac{\frac{\beta_{1}m_{{21}}p_{{21}}S_{2}}{m_{{11}}S_{1}+m_{{21}}S_{2}}+\frac{\beta_{2}m_{{22}}p_{{22}}S_{2}}{S_{1}m_{{12}}+S_{2}m_{{22}}}}{\gamma_{2}+\delta_{2}+\mu_{2}}\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}}$
The dominant eigenvalue for this particular system is an incredibly long algebraic expression. However, with the $F$ and $V$ matrices matching that of the system outlined in the paper [23] and subsequently for two patches and two groups in Equations 32-39, the resulting $R_{0}$ coming from the dominant eigenvalue will be equal.
4.7 SIR Vector-borne Model
The outline of the original ODE system coming from Wedajo et al. describes the SIR Malaria model [24], but the structure of this model can be taken more broadly to represent a simple example of vector-borne disease modeling. The SIR vector-borne model demonstrates the adaptability of our method to different modes of transmission, extending its applicability from directly transmitted diseases to complex vector-host systems. The compartments $S_{h},I_{h},R_{h},S_{v},$ and $I_{v}$ represent susceptible humans, infected humans, recovered humans, susceptible vectors, and infected vectors. Note that in a vector-borne model, the vector’s infected compartments are considered general infected compartments for purposes of calculating $R_{0}$. The vector for Malaria in the original model is mosquitoes. The parameters $\Pi,\beta_{hv},\mu_{h},\delta,\alpha,\sigma,\Lambda,\beta_{vh},$ and $\mu_{v}$ represent birth/immigration of humans, the rate of infected vectors infecting humans, human natural death rate, the rate of infected humans immigrating, the rate of infected humans emigrating, the recovery rate, the birth/immigration rate of the vector, the rate of infected humans infecting vectors, and natural vector death rate, respectively.
$\frac{dS_{h}}{dt}$ $=$ $\Pi-\beta_{hv}S_{h}I_{v}-\mu_{h}S_{h}$ (40)
$\frac{dI_{h}}{dt}$ $=$ $\beta_{hv}S_{h}I_{v}+\delta I_{h}...$ (41)
$-(\alpha+\sigma+\mu_{h})I_{h}$
$\frac{dR_{h}}{dt}$ $=$ $\sigma I_{h}-\mu_{h}R_{h}$ (42)
$\frac{dS_{v}}{dt}$ $=$ $\Lambda-\beta_{vh}S_{v}I_{h}-\mu_{v}S_{v}$ (43)
$\frac{dI_{v}}{dt}$ $=$ $\beta_{vh}S_{v}I_{h}-\mu_{v}I_{v}$ (44)
Figure 8: SIR Vector-borne Variable Arc Weight PN Model
First, we obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the non-infected places token values, $(S_{h}^{*},S_{v}^{*},I_{h}^{*},I_{v}^{*},R_{h}^{*})=(\frac{\Pi}{\mu_{h}},\frac{\Lambda}{\mu_{v}},0,0,0)$. Then the matrix $\mathcal{F}$ by looking at the arc weights of arcs going into the infected compartments giving us
$\mathcal{F}_{1,2}(x)$ $=$ $\begin{bmatrix}\beta_{hv}S_{h}I_{v}&\beta_{hv}S_{h}I_{v}\\ \beta_{vh}S_{v}I_{h}&\beta_{vh}S_{v}I_{h}\end{bmatrix}.$
From this, we find the rate of change with respect to $I_{h}$ in the first column and $I_{v}$ in the second column, and evaluate at the DFE to give
$F$ $=$ $\begin{bmatrix}0&\beta_{hv}\frac{\Pi}{\mu_{h}}\\ \beta_{vh}\frac{\Lambda}{\mu_{v}}&0\end{bmatrix}.$
For $\mathcal{V}$, we look at arc weights of arcs leaving the infected places and transitioning from one infected place to another, yielding
$\mathcal{V}_{1,2}(x)$ $=$ $\begin{bmatrix}\alpha I_{h}+\mu_{h}I_{h}+\sigma I_{h}-\delta I_{h}&0\\ 0&\mu_{v}I_{v}\end{bmatrix}.$
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this we find the inverse to give
$V$ $=$ $\begin{bmatrix}\alpha+\mu_{h}+\sigma-\delta&0\\ 0&\mu_{v}\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\alpha+\mu_{h}+\sigma-\delta}&0\\ 0&\frac{1}{\mu_{v}}\end{bmatrix}.$
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}F\cdot V^{-1}\end{bmatrix}=\sqrt{(\frac{\beta_{hv}\Pi}{\mu_{h}\mu_{v}}\cdot\frac{\beta_{vh}\Lambda}{(\mu_{v})(\alpha+\mu_{h}+\sigma-\delta)})}$
Some of the parameters are renamed in the lay out of the ODE model in equations 40-44 versus the original [24]. With this parameter renaming, the resulting $R_{0}$ value equals the exact $R_{0}$ expression in the Wedajo paper.
5 Numerical Verification
This section presents the numerical verification of the algebraic expression for the basic reproduction number, $R_{0}$, derived from our Petri net $R_{0}$ methodology laid out above. While the analytical derivation of $R_{0}$ from the structural properties of the Petri net provides a concise mathematical representation of $R_{0}$ using the VAPN method, this section aims to provide robust, simulation-based evidence to corroborate these theoretical findings. In fact, the numerical results presented offer an independent validation of our $R_{0}$ expression.
To achieve this validation, we conducted simulations of the Petri net model using GPenSim, a general-purpose Petri net simulator. GPenSim allows for close representation of ODE systems in Petri nets [7]. By analyzing the output of these simulations, specifically the progression of susceptible population and infectious population under various parameter sets, we can numerically estimate the $R_{0}$ exhibited by the simulated system. This analysis is independent of any algebraic $R_{0}$ methods that look at the parent system. The subsequent results will demonstrate the concordance between the $R_{0}$ value obtained through algebraic methods and the $R_{0}$ value inferred from the GPenSim simulation outcomes, thereby verifying the accuracy and reliability of our analytical expression. To show this we compare the $R_{0}$ values produced algebraically from the systems and through various methods of looking at the resulting simulation data. The comparison itself via relative root mean square error (RRMSE).
The Petri net models and ODE models have been defined and run for a range of parameters. For models SIRS 4.1 and the Nonlinear model 4.5, error surfaces were found for the $R_{0}$ values depending on three of the parameters. The $R_{0}$ values from the respective ODE system and PN system are found directly from the resulting data using R-studio package called R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics library from Boelle and Obadia [25, 26]. Both the Attack Rate method [1] and the Maximum Likelihood method [27] within the library were utilized. The $R_{0}$ values for each system are then compared across the range of parameters to the equivalent algebraically, NGM and NGM-PN $R_{0}$ value.
5.1 SIRS
The algebraic expression using the NGM and NGM-PN for the SIRS 4.1 systems is $R_{0}=\frac{\beta}{\gamma}$. The following Figure 9 compares the algebraic expression with that found from applying the attack rate method within the R library to the resulting simulation data across parameter values.
Petri Net $R_{0}$ vs Algebraic $R_{0}$ for SIRS system
Figure 9: Using the R library R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics from Boelle and Obadia [25, 26] with estimation method of $R_{0}$ of Attack rate. The error of the simulations remains below 1% RRMSE for all parameter values in the SIRS system and below 0.1% RRMSE for nearly all parameter values, including all biologically plausible values. Additionally, all $R_{0}$ values remain in the produced 95% confidence interval produced for the $R_{0}$ values given from the R library.
5.2 Nonlinear System
The algebraic expression of $R_{0}$ using the NGM and NGM-PN for the Nonlinear
refsec:NL systems is $R_{0}=R_{0}=\varrho(FV^{-1})=\frac{\sigma\beta}{(\sigma+\mu)(\gamma+\mu)}$. Since $R_{0}$ includes the parameters $\beta,\gamma,$ and $\sigma$ these will be used as the parameters where the $R_{0}$ values will be compared across a range for each.
Petri Net $R_{0}$ vs Algebraic $R_{0}$ for Nonlinear system
Figure 10: Using the R library R0: Estimation of R0 and Real-Time Reproduction Number from Epidemics from Boelle and Obadia [25, 26] with estimation method of $R_{0}$ of Attack rate. The error of the simulations remains below 1.2% RRMSE for all parameter values in the Nonlinear system and below 0.1% RRMSE for nearly all parameter values, including all biologically plausible values. Additionally, all $R_{0}$ values remain in the produced 95% confidence interval produced for the $R_{0}$ values given from the R library.
As shown in Figure 10, the Petri net implementation of the nonlinear model produces $R_{0}$ values that closely match the algebraic solution, with RRMSE consistently below 0.1%. This high level of agreement demonstrates the accuracy and numerical stability of the NGMPN method, even in the presence of nonlinear transmission and recovery terms. These results support the validity of the approach for more complex and mechanistically detailed epidemiological models.
6 Discussion and Conclusions
In this work, we devised a formal framework for finding the basic reproduction number $R_{0}$ within systems modeled by Petri nets. This framework draws upon well-established methodologies employed in the analysis of ODE models. The relevance of applying $R_{0}$ in the context of Petri nets is particularly pronounced, given the increasing utilization of Petri nets in epidemiological modeling to depict more intricate and dynamic scenarios of disease spread. These scenarios include multi-compartment models, spatially distributed populations, and frameworks incorporating social and nonlinear dynamics. Our approach builds on the next-generation matrix method, which is well understood within the realm of ODEs. Our results extend the definition of $R_{0}$ to Petri nets through the interpretation of token transitions and arc weights in a manner that aligns with underlying epidemiological processes.
We demonstrate that the NGM methodology can be effectively employed in the context of Petri nets by first defining relevant matrices that represent the rate of infection $F$ and the rate of removal or transition out of compartments $V$. Subsequently, we compute the dominant eigenvalue of the product of $FV^{-1}$. This approach is consistent with established NGM techniques for ODE frameworks, resulting in a robust mechanism for quantifying the basic reproduction number when applied to Petri nets. The examples elucidated—including models such as SIRS, SEEIR, Basic COVID-19, Nonlinear, Patch, and SIR Vector-borne serve to illustrate the versatility of this methodology across a spectrum of disease dynamics characterized by various place to transition structures and arc weights.
A primary challenge encountered in adapting the NGM framework to Petri nets pertains to effectively translating the compartmental structures and transition rates—typically articulated through arc weights and token movement—into a form compatible with the mathematical formulations employed in the NGM. Nevertheless, once the transition and removal rates are contextualized in terms of Petri net parameters using Diekmann’s approach [5], the ensuing process of calculating $R_{0}$ becomes straightforward and closely parallels that utilized for ODEs. This provides a powerful and generalizable method for scrutinizing disease dynamics in models based on Petri nets.
Moreover, the framework afforded by Petri nets enables the integration of more complex features that are often challenging to model using traditional ODE approaches. These features include variable parameters, discrete events, and non-linear transitions [9]. Such flexibility renders Petri nets a compelling tool for accurately representing real-world epidemiological systems that exhibit complex interactions and dependencies, including co-infections, varying transmission rates, and spatial heterogeneity in disease dissemination.
The proposed method, maintains practical limitations that also limit other methods of $R_{0}$ including those used for ODEs. These practical challenges persist in its application to extensive or highly intricate Petri nets. Specifically, the computational demands associated with calculating the dominant eigenvalue of large matrices can be substantial, particularly when addressing large-scale disease models that encapsulate numerous compartments. Again, these problems currently exist with all methods of algebraic $R_{0}$ formulation, but should be noted. Additionally, considerations pertaining to the discrete nature of token movement in Petri nets may necessitate careful deliberation when modeling diseases characterized by continuous dynamics.
Future inquiries may be directed towards enhancing computational techniques to accommodate larger and more intricate systems, as well as exploring the application of this method to empirical epidemiological data. Overall, this work contributes to bridging the divide between traditional ODE-based modeling approaches and Petri net frameworks, thereby providing a more comprehensive toolkit for understanding and predicting disease propagation across a variety of contexts.
Funding
Trevor Reckell is partially supported by NIH grant DMS-1615879. Petar Jevtić and Beckett Sterner are supported by NIH grant 5R01GM131405-02.
Conflict of Interest
The authors declare no conflict of interest.